Todrick Company is a merchandiser that reported the following information based on 1,000 units sold:

Sales $ 465,000
Beginning merchandise inventory $ 31,000
Purchases $ 310,000
Ending merchandise inventory $ 15,500
Fixed selling expense $ ?
Fixed administrative expense $ 18,600
Variable selling expense $ 23,250
Variable administrative expense $ ?
Contribution margin $ 93,000
Net operating income $ 27,900

Required:

1. Prepare a contribution format income statement.

2. Prepare a traditional format income statement.

3. Calculate the selling price per unit.

4. Calculate the variable cost per unit.

5. Calculate the contribution margin per unit.

6. Which income statement format (traditional format or contribution format) would be more useful to managers in estimating how net operating income will change in responses to changes in unit sales?
Answer #1

Answer:

Cost of Goods Sold = Beginning Inventory + Purchases - Ending Inventory
Cost of Goods Sold = $31,000 + $310,000 - $15,500
Cost of Goods Sold = $325,500

Contribution Margin = Sales - Variable Expenses
$93,000 = $465,000 - Variable Expenses
Variable Expenses = $372,000

Variable Expenses = Cost of Goods Sold + Variable Selling Expense + Variable Administrative Expenses
$372,000 = $325,500 + $23,250 + Variable Administrative Expenses
Variable Administrative Expenses = $23,250

Net Operating Income = Contribution Margin - Fixed Expenses
$27,900 = $93,000 - Fixed Expenses
Fixed Expenses = $65,100

Fixed Expenses = Fixed Selling Expenses + Fixed Administrative Expenses
$65,100 = Fixed Selling Expense + $18,600
Fixed Selling Expenses = $46,500

Answer to Part 3.

Selling price per Unit = Total Sales / Units Sold
Selling price per Unit = 465,000 / 1,000
Selling price per Unit = $465

Answer to Part 4.

Variable Cost per Unit = Total Variable Expense / Units Sold
Variable Expenses = Cost of Goods Sold + Variable Selling Expense + Variable Administrative Expenses
Variable Expense = $325,500 + $23,250 + $23,250
Variable Expenses = $372,000

Variable Cost per Unit = 372,000 / 1,000
Variable Cost per Unit = $372

Answer to Part 5.

Contribution Margin per Unit = Contribution Margin / Units Sold
Contribution Margin per Unit = 93,000 / 1,000
Contribution Margin per Unit = $93